OpenStudy (anonymous):

passing through (4,-2) and perpendicular to x=5/4y-2. Write in slope intercept form.

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

is the negative 2 also in the denominator? It's a little hard to read correctly without brackets.\[\large x=\frac{5}{4y-2}\]

zepdrix (zepdrix):

\[\large x=\frac{5}{4}y-2\]Oh like this? :) That would make more sense lol

OpenStudy (anonymous):

no it isn't. it's like the second one you just put up.

zepdrix (zepdrix):

Let's first get our line into slope-intercept form, so we can accurately identify the slope. We want it in this form,\[\large y=mx+b\] \[\large x=\frac{5}{4}y-2\]Start by adding 2 to each side,\[\large x+2=\frac{5}{4}y\]Multiply each side by 4/5,\[\large \frac{4}{5}(x+2)=\left(\frac{5}{4}y\right)\frac{4}{5}\]The fractions will cancel out on the right, giving us,\[\large y=\frac{4}{5}x+\frac{8}{5}\] So we've got our line in slope-intercept form, this will make it easier to work with. Understand those steps so far?

OpenStudy (anonymous):

oh i had put a -4/5

OpenStudy (anonymous):

but yes other than that i get it so far

zepdrix (zepdrix):

So it looks like our slope \(\large m\) is \(\large \dfrac{4}{5}\). Do you know what it means for a line to be `perpendicular`? It relates to the slope. :) Since the line we're trying to form is perpendicular to this line, it will have a slope that is a `negative reciprocal` of this slope. Do you understand what that means? c:

OpenStudy (anonymous):

-4/5 correct? that would be the reciprocal

zepdrix (zepdrix):

Hmm I believe it's going to be, -5/4. Yes the negative looks good. But then we take the reciprocal (the flip) of our fraction.

OpenStudy (anonymous):

oh okay

zepdrix (zepdrix):

So we're trying to form an equation for a line perpendicular to the given line. Let's call this new line something likeeeee \(\large y_p\). So we're trying to get an equation for this line, \(\large y_p=mx+b\). We've determined that the slope is, \(\large m=-\dfrac{5}{4}\). Now we need to find the \(\large b\) value, the `y-intercept`. To do so, we'll plug in the coordinate pair they gave us, that this line passes through.

zepdrix (zepdrix):

So plug \(\large (4,-2)\) into our equation \(\large y_p=-\dfrac{5}{4}x+b\).

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i think im messing up so where

OpenStudy (anonymous):

for som reason im getting y=-4/5x+6/5

OpenStudy (anonymous):

i mean -5/4x+6/5

zepdrix (zepdrix):

So plugging in our point gives us,\[\large -2=-\frac{5}{4}(4)+b\]The 4's will cancel out,\[\large -2=-\frac{5}{\cancel4}(\cancel4)+b\]Giving us,\[\large -2=-5+b\]Adding 5 to each side,\[\large b=3\] Did you do something different?

OpenStudy (anonymous):

yeah i did

OpenStudy (anonymous):

i was doing it wrong

OpenStudy (anonymous):

so y=4/5x+8/5 is correct?

zepdrix (zepdrix):

This was the equation of the line we started with, \[\large y=\frac{4}{5}x+\frac{8}{5}\] They gave us a bunch of information to find a different line. And that was this line,\[\large y_p=mx+b\] We determined that the slope is \(\large m=-\dfrac{5}{4}\). And that the y-intercept is \(\large b=3\).

zepdrix (zepdrix):

The p doesn't mean anything, you don't need to put that if you don't want. It was just so we could tell it apart from our original y, which referred to a different line.

OpenStudy (anonymous):

so the answer is y=-5/4x+3

zepdrix (zepdrix):

ya :)